OpenStudy (anonymous):

I WILL FAN Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (anonymous):

@anthonyc123 Hello??? Are you actually typing?

OpenStudy (anonymous):

Do see which linear equation that this point works for, simply plus in the X and Y values for that point into each equation, and see if it holds true. First of all, they are all in slope-intercept form. \[y=mx +b\] where m=slope. Since the slope needs to be -3, and all of these have the slope, you can't rule any of them out this way. Option 1: \[y =-3x -13\] with point (2,-5) for (x,y) \[(-5)=-3(2)-13\] \[-5=-6-13 ?\] \[-5=-19 ?\] NO, this point is not a value for this linear equation. Option 2: \[y =-3x +11\] with point (2,-5) for (x,y) \[(-5)=-3(2)+11\] \[-5=-6+11 ?\] \[-5=5\] NO, this point is not a value for this linear equation. Option 3: \[y =-3x +13\] with point (2,-5) for (x,y) \[(-5)=-3(2)+13 ?\] \[-5=-6+13 ?\] \[-5=7 ?\] NO, this point is not a value for this linear equation. Option 4: \[y =-3x +1\] with point (2,-5) for (x,y) \[(-5)=3(2)+1 ?\] \[-5=-6+1 ?\] \[-5=-5 \] YES, this point is in this linear equation.

OpenStudy (anonymous):

@anthonyc123 Thank you so much!!!!!

OpenStudy (anonymous):

MMhmm